verify maximum Q =( L ,K) subjected to c=wage×labour +interest×capital
1
Expert's answer
2021-08-25T17:44:15-0400

Solution:

The cost-minimizing combination of capital and labor is the one where:


"MRTS_{K,L} = \frac{ MP_{L} }{ MP_{K} } = \frac{ w}{ r }"

C = WL + rK


"\frac{ MP_{L} }{ w } = \frac{ MP_{K} }{ r } = \frac{ MP_{L} }{ MP_{K} } = \frac{ w }{ r}"


Therefore:


"MRTS_{K,L} = \frac{ w}{ r }"